A mixture of 0.100 mol of NO, 0.050 mol of H2 and 0.10 mol of H2O is placed in a 1.0-L vessel at 300 K. The following equilibrium is established:

2NO(g) + 2H2(g) ⇆ N2(g) + 2H2O(g).


At equilibrium [NO] = 0.062M. Calculate the equilibrium concentrations of H2, N2 and H2O and Keq.


1
Expert's answer
2022-06-07T09:05:03-0400

Initially,

[NO]= 0,1mol/l; [H2]= 0,05mol/l; [H2O]= 0,1mol/l

After chemical equilibrium [NO]= 0,062mol/l

we find how many reacts NO:

0,1-0,062 = 0,038

2NO + 2H2 = N2 + 2H2O

0,038 — 0,038 — 0,019 — 0,038

[H2]equ = 0,05-0,038 = 0,012 mol/l

[N2]equ = 0,019 mol/l

[H2O]equ = 0,1+0,038 = 1,038 mol/l

[NO]equ = 0,062 mol/l


"K_{equ}= \frac {[N_2][H_2O]}{[NO][H_2]}= \frac{0,019*1,038}{0,062*0,012}= 26,5"


Answer:

[H2]equ = 0,012 mol/l

[N2]equ = 0,019 mol/l

[H2O]equ = 1,038 mol/l

Kequ = 26,5